Help with this problem Please. Find the eqaution of each line. Write the eqaution in standard form unless otherwise.
Slope 2; through (-2,3)

Answer by Alan3354(69443)

Slope 2; through (-2,3)
y-y1 = m*(x-x1) the point is (x1,y1)
y-3 = 2(x+2)
y-3 = 2x+4
2x-y = -7